# GDBA Revisited: Unleashing the Power of Guided Local Search for Distributed Constraint Optimization

### v1

@@ 4 Empirical Evaluations
We evaluate our algorithm on various standard DCOP benchmarks, including: (1) random DCOPs with 120 agents, a graph density of 0.1 (sparse) or 0.6 (dense); (2) scale-free networks (Barabási and Albert, 1999) with 120 agents and $m_{0}=m_{1}=3$; (3) 2D lattices with grid size of $10\times 10$; (4) meeting scheduling problems using Events-as-Variable (EAV) formulation (Zivan et al., 2014; Maheswaran et al., 2004b) with 20 available time slots, 20 meetings and 90 persons, where each person randomly selects two meetings to participate, the travel time between any pair of meetings is uniformly drawn from $[6,10]$, and a cost equal to the number of overbooked persons is incurred if the difference between the time slots of two meetings are less than the travel time; (5) weighted graph coloring problems with 120 agents, 3 available colors, and graph density of 0.05, where a conflict cost is uniformly selected from $[1,100]$ if two adjacent agents assign the same color. For benchmarks (1–3), we consider the problems with a domain size of 10 and uniformly select constraint costs from $[0,100]$. For each benchmark, we generate 100 random instances and each instance is solved 20 times with the maximum round of 1000. Finally, we average the anytime cost (Zilberstein, 1996; Zivan et al., 2014) of each round of 2000 runs as the final result. All experiments are conducted on a Linux workstation with Intel Xeon W-2133 CPU and 32GB memory.
For competitors, we consider DSA (Zhang et al., 2005) with $p=0.8$ and GDBA (Okamoto et al., 2016) as representative local search algorithms; MGM2 (Maheswaran et al., 2004a) as a representative $K$-OPT algorithm, and Damped Max-sum (Cohen et al., 2020) with damping factors of 0.7 and 0.9 as the state-of-the-art baselines. For GDBA, we consider its $(M,NM,T)$ variant that exhibits the strongest performance according to (Okamoto et al., 2016).

### v2

@@ 4 Empirical Evaluations
We evaluate our algorithm on various standard DCOP benchmarks, including: (1) random DCOPs with 120 agents, a graph density of 0.1 (sparse) or 0.6 (dense); (2) scale-free networks (Barabási and Albert 1999) with 120 agents and $m_{0}=m_{1}=3$; (3) 2D lattices with grid size of $10\times 10$; (4) meeting scheduling problems using Events-as-Variable (EAV) formulation (Zivan et al. 2014; Maheswaran et al. 2004b) with 20 available time slots, 20 meetings and 90 persons, where each person randomly selects two meetings to participate, the travel time between any pair of meetings is uniformly drawn from $[6,10]$, and a cost equal to the number of overbooked persons is incurred if the difference between the time slots of two meetings are less than the travel time; (5) weighted graph coloring problems with 120 agents, 3 available colors, and graph density of 0.05, where a conflict cost is uniformly selected from $[1,100]$ if two adjacent agents assign the same color. For benchmarks (1–3), we consider the problems with a domain size of 10 and uniformly select constraint costs from $[0,100]$. For each benchmark, we generate 100 random instances and each instance is solved 20 times with the maximum round of 1000. Finally, we average the anytime cost (Zilberstein 1996; Zivan et al. 2014) of each round of 2000 runs as the final result. All experiments are conducted on a Linux workstation with Intel Xeon W-2133 CPU and 32GB memory.
For competitors, we consider DSA (Zhang et al. 2005) with $p=0.8$ and GDBA (Okamoto et al. 2016) as representative local search algorithms; MGM2 (Maheswaran et al. 2004a) as a representative $K$-OPT algorithm, and DMS (Cohen et al. 2020) with damping factors $\lambda=0.7$ and 0.9 as the state-of-the-art baselines. For GDBA, we consider its $(M,NM,T)$ variant that exhibits the strongest performance according to (Okamoto et al. 2016).
@@ 5 Conclusion
Acknowledgments
This research is supported by the Ministry of Education, Singapore, under its MOE AcRF Tier 2 Award MOE-T2EP20223-0003. Xinrun Wang is supported by Singapore Ministry of Education (MOE) Academic Research Fund (AcRF) Tier 1 grant (No. MSS24C005). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.